# Relative portfolio optimization via a value at risk based constraint

### v1

@@ 6. Numerical Examples
Let us discuss the Nash equilibria from Theorems 3.3, 3.5 and 4.2 by considering some numerical examples. To do this, we use a Black-Scholes financial market consisting of one stock with price process
@@
with $\nu=-0.0225$ and $\tau^{2}=0.045.$
First, we consider the 2-agent equilibrium under logarithmic utilities from Theorem 3.3. We choose $x_{0}^{1}=3$ and $x_{0}^{2}=2$. Figure 1 shows the Nash equilibrium $(X_{1}^{*},X_{2}^{*})$ as a function of the state-price density for $\alpha_{2}=0.2$ and $\beta_{1}=0.9$, i.e. $\alpha_{2}\beta_{1}x_{0}^{1}\leq x_{0}^{2}<\beta_{1}x_{0}^{1}$, so that the Nash equilibrium is given in the second part of Theorem 3.3 c). The set $A_{2}$ is chosen as $A_{2}=\{Z_{T}\leq z_{\alpha_{2}}\}$ as discussed in Remark 3.4. The purple and green solid lines show the terminal wealth of agent 1 and 2 in the Nash equilibrium. The blue dashed line shows for comparison the optimal terminal wealth of agent 2 in the standard Merton problem without the VaR-based constraint. The terminal wealth $X_{1}^{*}$ is continuous, strictly decreasing, and strictly convex in terms of $Z_{T}$ while $X_{2}^{*}$ shows a similar overall behavior with a discontinuity located at $z_{\alpha_{2}}$. We notice, that the terminal wealth $X_{2}^{*}$ in the Nash equilibrium is larger than the standard solution if $Z_{T}\leq z_{\alpha_{2}}$ and smaller for $Z_{T}>z_{\alpha_{2}}$.
@@
Figure 2(b) shows the terminal wealth $X_{2}^{*}$ of agent 2 in the Nash equilibrium for different choices of the parameter $\beta_{1}$. We notice a change of order of the value of $X_{2}^{*}$ for the different choices of $\beta_{1}$ located at the discontinuity $z_{\alpha_{2}}.$ For $Z_{T}\leq z_{\alpha_{2}}$, the value of $X_{2}^{*}$ is largest for the largest choice of $\beta_{1}$ as the value of $X_{2}^{*}$ is kept at $\beta_{1}X_{1}^{*}$ in this case. For $Z_{T}>z_{\alpha_{2}}$, the order of the values changes, i.e. the largest choice of $\beta_{1}$ yields the smallest value of $X_{2}^{*}$. This is again due to the budget constraint $\mathbb{E}[Z_{T}X_{2}^{*}]=x_{0}^{2}$.
Finally, let us consider some different possible choices for the set $A_{2}$. As discussed in Remark 3.4, there are infinitely many possible choices for $A_{2}$. Although choosing $A_{2}=\{Z_{T}\leq z_{\alpha_{2}}\}$ maximizes the expected terminal wealth, it is worth considering different choices for $A_{2}$ and their influence on the terminal wealth $X_{2}^{*}$ of agent 2 in the Nash equilibrium. Figure 3 shows $X_{2}^{*}$ for different choices of $A_{2}$ in the form $A_{2}=\{b\leq Z_{T}\leq c\}$. Since $Z_{T}\sim\mathrm{LN}(\nu,\tau)$, for a fixed lower bound $b>0$, the upper bound $c>0$ of the interval is determined via
$c=\exp\bigg{(}\nu+\tau\cdot\Phi^{-1}\bigg{(}\alpha_{2}+\Phi\bigg{(}\frac{\mathrm{ln}(b)-\nu}{\tau}\bigg{)}\bigg{)}\bigg{)}.$
Note that the largest possible choice for the lower bound $b$ is the $(1-\alpha_{2})$-quantile of $Z_{T}$, i.e.
@@
where $\Phi$ denotes the cumulative distribution function of the standard normal distribution. Figures 4, 5(a) and 5(b) illustrate the Nash equilibrium in terms of $Z_{T}$ for the parameter choice $x_{0}^{1}=3$, $x_{0}^{2}=2$, $\gamma=0.7$ and $\alpha_{i},\beta_{i}$, $i=1,2,$ so that (3.19) holds.
Figure 4 shows the Nash equilibrium $(X_{1}^{*},X_{2}^{*})$ as a function of $Z_{T}$ for $\alpha_{2}=0.5$ and $\beta_{1}=0.9$. The purple and blue solid line show the terminal wealth of agents 1 and 2 in the Nash equilibrium. For comparison, we also illustrated the unconstrained optimal terminal wealth of agent 2 as the green dashed line. Similar to the logarithmic case, the terminal wealth of agent 1 is again the solution to the standard problem without the VaR-based constraint and is thus continuous. The terminal wealth $X_{2}^{*}$ of agent 2 is smaller than the unconstrained terminal wealth for $Z_{T}<z_{1-\alpha_{2}}$ and larger for $Z_{T}\geq z_{1-\alpha_{2}}$ due to the structure of $X_{2}^{*}$ shown in Theorem 3.5.
Figure 5(a) illustrates the influence of the parameter $\alpha_{2}$ on $X_{2}^{*}$. As expected, the location of the discontinuity of $X_{2}^{*}$ is decreasing in terms of $\alpha_{2}$ as it is simply given as the $(1-\alpha_{2})$-quantile of $Z_{T}$. Moreover, we notice that for $Z_{T}<z_{\alpha_{2}}$ the value of $X_{2}^{*}$ is decreasing in $\alpha_{2}$. This results from the budget constraint $\mathbb{E}[Z_{T}X_{2}^{*}]=x_{0}^{2}$ since for larger $\alpha_{2}$, the value of $X_{2}^{*}$ is kept on the larger value $\beta_{1}X_{1}^{*}$ on a larger interval.
Finally, Figure 5(b) shows the terminal wealth $X_{2}^{*}$ of agent 2 in the Nash equilibrium for different choices of the parameter $\beta_{1}$. We notice a change of order of the value of $X_{2}^{*}$ for the different choices of $\beta_{1}$ located at the discontinuity $z_{1-\alpha_{2}}.$ For $Z_{T}\geq z_{1-\alpha_{2}}$, the value of $X_{2}^{*}$ is largest for the largest choice of $\beta_{1}$ as the value of $X_{2}^{*}$ is kept at $\beta_{1}X_{1}^{*}$ in this case. For $Z_{T}<z_{1-\alpha_{2}}$, the order of the values is opposite, i.e. the largest choice of $\beta_{1}$ yields the smallest value of $X_{2}^{*}$. This is again due to the budget constraint $\mathbb{E}[Z_{T}X_{2}^{*}]=x_{0}^{2}$.
@@
It remains to discuss Nash equilibria for more than two agents using logarithmic utilities. Here, we restrict to the case that the sum $\alpha_{1}+\ldots+\alpha_{n}$ is less or equal to $1$. Thus, we can use Theorem 4.2 to compute the terminal wealth of the agents in the Nash equilibrium. Figure 6 displays the terminal wealths of $n=4$ agents in the Nash equilibrium for the parameter choice $\alpha_{i}=0.2$, $\beta_{i}=0.3$ for all $i\in\{1,\ldots,4\}$ and $x_{0}^{1}=5$, $x_{0}^{2}=4$, $x_{0}^{3}=3$, $x_{0}^{4}=2$. The sets $A_{i}$, $i=1,\ldots,4,$ are chosen as $A_{i}=(z_{\alpha_{1}+\ldots+\alpha_{i-1}},z_{\alpha_{1}+\ldots+\alpha_{i}}]$, $i=1,\ldots,4$, where $z_{\alpha_{1}+\ldots+\alpha_{i}}:=0$ for $i=0$. We notice that the terminal wealth of agent 4 is larger than the terminal wealth of agent 3 on the set $A_{4}$ although agent 4 starts with a smaller initial capital. However, this results in a terminal wealth on $A_{4}^{c}$ that is significantly smaller than the optimal terminal wealth in the respective unconstrained problems. A comparison of $X_{3}^{*}$ and $X_{4}^{*}$ to the solutions of the respective unconstrained problems can be found in Figures 7(a) and 7(b).
Figure 6. Nash equilibrium $(X_{1}^{*},X_{2}^{*},X_{3}^{*},X_{4}^{*})$ from Theorem 4.2 in terms of $Z_{T}$ for $\alpha_{i}=0.2$, $\beta_{i}=0.3$ for $i\in\{1,\ldots,4\}$, $x_{0}^{1}=5$, $x_{0}^{2}=4$, $x_{0}^{3}=3$, $x_{0}^{4}=2$ and $A_{i}=(z_{\alpha_{1}+\ldots+\alpha_{i-1}},z_{\alpha_{1}+\ldots+\alpha_{i}}]$, $i=1,\ldots,4$, where $z_{\alpha_{1}+\ldots+\alpha_{i}}:=0$ for $i=0$.
@@ 7. Conclusion
It is meaningful to consider competing agents by introducing probability constraints instead of relative wealth targets. However, it turns out that finding Nash equilibria is in general more challenging. An interesting case which is analytically solvable is the case that all investors have a logarithmic utility function. In this case, Nash equilibria (if they exist) can be found for an arbitrary number of agents. In order to succeed, a number of parameter cases have to be distinguished leading either to no Nash equilibrium, a unique Nash equilibrium or an infinite number of Nash equilibria. We illustrate our findings by numerical examples. For other utility functions, and in particular, if the agents have different utility functions, the problem is much more demanding and we leave it for future research.
@@ 8.3. Maximizing $L$
$L(X,\lambda_{2},\eta_{2}):=U(X)-\lambda_{2}Z_{T}X+\eta_{2}\mathbbm{1}_{[X\geq\beta_{1}I(\lambda_{1}Z_{T})]}$
with $\lambda_{2},\eta_{2}\geq 0$ fixed. We show that $L$ is maximized by $X_{2}^{*}$ from (3.20). Obviously the maximum points can either be $I(\lambda_{2}Z_{T})$ or $\beta_{1}I(\lambda_{1}Z_{T})$ where $I(\lambda_{2}Z_{T})<\beta_{1}I(\lambda_{1}Z_{T}).$ We can compare the two possible values of the function $L$:
$F(Z_{T}):=U(I(\lambda_{2}Z_{T}))-\lambda_{2}Z_{T}I(\lambda_{2}Z_{T})-\Big{\{}U(\beta_{1}I(\lambda_{1}Z_{T}))-\lambda_{2}Z_{T}\beta_{1}I(\lambda_{1}Z_{T})+\eta_{2}\Big{\}}.$
@@
$F^{\prime}(Z_{T})$ $=\lambda_{2}\big{(}\beta_{1}I(\lambda_{1}Z_{T})-I(\lambda_{2}Z_{T})\big{)}-\beta_{1}\lambda_{1}Z_{T}I^{\prime}(\lambda_{1}Z_{T})(\lambda_{1}\beta_{1}^{-\gamma}-\lambda_{2})$
$=Z_{T}^{-1/\gamma}\bigg{(}\lambda_{2}\Big{(}\beta_{1}\lambda_{1}^{-1/\gamma}-\lambda_{2}^{-1/\gamma}\Big{)}+\frac{\beta_{1}\lambda_{1}^{-1/\gamma}}{\gamma}\Big{(}\lambda_{1}\beta_{1}^{-\gamma}-\lambda_{2}\Big{)}\bigg{)}=:Z_{T}^{-1/\gamma}f(\lambda_{2}).$
@@
$f\big{(}\lambda_{1}\beta_{1}^{-\gamma}\big{)}=0.$
@@
$f^{\prime}(\lambda_{2})$ $=\Big{(}1-1/\gamma\Big{)}\Big{(}\beta_{1}\lambda_{1}^{-1/\gamma}-\lambda_{2}^{-1/\gamma}\Big{)},$
so that $f^{\prime}(\lambda_{2})<0$ for all $\lambda_{2}>\lambda_{1}\beta_{1}^{-\gamma}$. Thus, since we already saw in the proof of Theorem 3.5 that $\lambda_{2}>\lambda_{1}\beta_{1}^{-\gamma}$, we deduce that $F^{\prime}(Z_{T})<0$ for all values of $Z_{T}$. Moreover, by definition of $\eta_{2}$, we have $F(z_{1-\alpha_{2}})=0$ which implies that $X_{2}^{*}$ maximizes the function $L$.

### v2

@@ 5. Numerical Examples
Let us discuss the Nash equilibria from Theorems 3.3, 3.6 and 4.2 by considering some numerical examples. To do this, we use a Black-Scholes financial market consisting of one stock with price process
@@
with $\nu=-0.045$ and $\tau^{2}=0.09.$ It is important to note here that large $Z_{T}$ correspond to small stock prices and vice versa.
5.1. Two agent case with logarithmic utility
First, we consider the 2-agent equilibrium under logarithmic utilities from Theorem 3.3. We choose $x_{0}^{1}=3$ and $x_{0}^{2}=2$. Figure 1 shows the Nash equilibrium $(X_{1}^{*},X_{2}^{*})$ as a function of the state-price density for $\alpha_{2}=0.2$ and $\beta_{1}=0.9$, i.e. $\alpha_{2}\beta_{1}x_{0}^{1}\leq x_{0}^{2}<\beta_{1}x_{0}^{1}$, so that the Nash equilibrium is given in the second part of Theorem 3.3 c). Note that $\alpha_{1}$ and $\beta_{2}$ have no influence on the Nash equilibrium. The set $A_{2}$ is chosen as $A_{2}=\{Z_{T}\leq z_{\alpha_{2}}\}$ as discussed in Remark 3.4. The purple and green solid lines show the terminal wealth of agent 1 and 2 in the Nash equilibrium. The black dashed line shows for comparison the optimal terminal wealth of agent 2 in the standard Merton problem without the VaR-based constraint. The terminal wealth $X_{1}^{*}$ is continuous, strictly decreasing, and strictly convex in terms of $Z_{T}$ while $X_{2}^{*}$ shows a similar overall behavior with a discontinuity located at $z_{\alpha_{2}}$. We notice, that the terminal wealth $X_{2}^{*}$ in the Nash equilibrium is larger than the standard solution if $Z_{T}\leq z_{\alpha_{2}}$ and smaller for $Z_{T}>z_{\alpha_{2}}$. This means that the agent insures the constraint in scenarios with a bullish stock price evolution and underperforms in bearish markets. However, as discussed before, the insured scenarios can to some extend be arbitrarily defined in the case of an agent with log-utility, as long as the scenario set has the desired probability.
@@ 5.1. Two agent case with logarithmic utility
Figure 2(b) shows the terminal wealth $X_{2}^{*}$ of agent 2 in the Nash equilibrium for different choices of the parameter $\beta_{1}$. We notice a change of order of the value of $X_{2}^{*}$ for the different choices of $\beta_{1}$ located at the discontinuity $z_{\alpha_{2}}.$ For $Z_{T}\leq z_{\alpha_{2}}$, the value of $X_{2}^{*}$ is largest for the largest choice of $\beta_{1}$ as the value of $X_{2}^{*}$ is kept at $\beta_{1}X_{1}^{*}$ in this case. For $Z_{T}>z_{\alpha_{2}}$, the order of the values changes, i.e. the largest choice of $\beta_{1}$ yields the smallest value of $X_{2}^{*}$. This is again due to the budget constraint $\mathbb{E}[Z_{T}X_{2}^{*}]=x_{0}^{2}$. It is interesting to see here that the set of insured scenarios is not influenced by $\beta_{1}.$ This parameter only has an influence on the severity of underperformance in non-insured states.
Next, let us consider some different possible choices for the set $A_{2}$. As discussed in Remark 3.4, there are infinitely many possible choices for $A_{2}$. Although choosing $A_{2}=\{Z_{T}\leq z_{\alpha_{2}}\}$ maximizes the expected terminal wealth, it is worth considering different choices for $A_{2}$ and their influence on the terminal wealth $X_{2}^{*}$ of agent 2 in the Nash equilibrium. Figure 3 shows $X_{2}^{*}$ for different choices of $A_{2}$ in the form $A_{2}=\{c_{1}\leq Z_{T}\leq c_{2}\}$. Since $Z_{T}\sim\mathrm{LN}(\nu,\tau)$, for a fixed lower bound $c_{1}>0$, the upper bound $c_{2}>0$ of the interval is determined via
$c_{2}=\exp\bigg{(}\nu+\tau\cdot\Phi^{-1}\bigg{(}\alpha_{2}+\Phi\bigg{(}\frac{\mathrm{ln}(c_{1})-\nu}{\tau}\bigg{)}\bigg{)}\bigg{)}.$ (5.3)
Note that the largest possible choice for the lower bound $c_{1}$ is the $(1-\alpha_{2})$-quantile of $Z_{T}$, i.e.
@@
Finally, let us take a look at the replicating strategies determined in Theorem 3.5. Figure 4 shows a comparison of one path (i.e. for one realization of the path $(Z_{t}),\,t\in[0,T],$ of the state price density) of the replicating strategies in terms of invested amounts in the benchmark case without the VaR-based constraint and for the Nash equilibrium. We displayed three different choices of the set $A$ from Theorem 3.5. Each set is of the form $A=\{c_{1}<Z_{T}<c_{2}\}$, where $c_{1}$ takes values in $\{0,0.75,1.2\}$ and $c_{2}$ is defined as in (5.3). We notice that the highest fluctuation of the amount invested into the stock appears towards the end of the time interval. This is a result from the VaR-based constraint which compares the wealth of agent 2 to the wealth of her competitor at the terminal time $T$. Thus, at the beginning of the investment period, the dominating investment motive is the own utility. Moreover, we notice that the investment behavior associated to $c_{1}=0$ and $c_{1}=0.75$ is a lot riskier in the depicted scenario than the benchmark portfolio towards the end of the time interval, while $c_{1}=1.2$ appears to be less risky as the amount invested into the stock is always smaller than the benchmark portfolio. However, the behavior heavily depends on the specific path evolution. See also Figure 6 which displays five different realizations of the replicating portfolio processes for $X_{2}^{*}$ from Theorem 3.3 c). For each path, we used $c_{1}=0$, i.e. the insured set is $A=\{Z_{T}\leq z_{\alpha_{2}}\}$. The plot shows a strong difference between the different realizations, especially when comparing the first, second, and fifth path. In general however, one can conclude that the probability constraint becomes important towards the end of the trading horizon and leads at least in some cases to risky and very volatile investment strategies. These are scenarios where towards the end we have a close competition between agents. Figure 5 displays the wealth processes associated to the strategies from Figure 4. The processes behave quite similar for the first part of the time interval. The largest difference can again be observed towards the terminal time $T$. Finally, Figure 7 shows the wealth processes to the strategies shown in Figure 6. The target terminal wealth $\beta_{1}X_{1}^{*}$ is shown as a square in the respective color at the terminal time. We notice that the target is achieved in one of the five realizations which aligns with the parameter choice of $\alpha_{2}=0.2$ in this case. Moreover, it is interesting to see that the simulation confirms that the insured set consists of scenarios where we have a good market performance and the target is hit precisely (of course this is only possible here since we assume no market frictions). In those scenarios where the target is not met, we can see that agents are quite far away from the target and underperform significantly.
Figure 4. Replicating strategy $\pi_{2}^{*}(t),$ $t\in[0,T]$, for the terminal wealth $X_{2}^{*}$ in Theorem 3.3 c) for $c_{1}\in\{0,0.75,1.2\}$ and the model parameters $\mu=0.03,\,\sigma=0.2,\,T=4,\,\alpha_{2}=0.2,\,\beta_{1}=0.9,\,x_{0}^{1}=3,\,x_{0}^{2}=2$.
Figure 5. Wealth process $X_{2}^{*}(t),\,t\in[0,T],$ for the terminal wealth $X_{2}^{*}$ in Theorem 3.3 c) for $c_{1}\in\{0,0.75,1.2\}$ and the model parameters $\mu=0.03,\,\sigma=0.2,\,T=4,\,\alpha_{2}=0.2,\,\beta_{1}=0.9,\,x_{0}^{1}=3,\,x_{0}^{2}=2$.
Figure 6. Five realizations of the replicating strategy $\pi_{2}^{*}(t),$ $t\in[0,T]$, for the terminal wealth $X_{2}^{*}$ in Theorem 3.3 c) for $c_{1}=0$ and the model parameters $\mu=0.03,\,\sigma=0.2,\,T=4,\,\alpha_{2}=0.2,\,\beta_{1}=0.9,\,x_{0}^{1}=3,\,x_{0}^{2}=2$.
Figure 7. Five realizations of the wealth process $X_{2}^{*}(t),$ $t\in[0,T]$, associated to the terminal wealth $X_{2}^{*}$ in Theorem 3.3 c) for $c_{1}=0$ and the model parameters $\mu=0.03,\,\sigma=0.2,\,T=4,\,\alpha_{2}=0.2,\,\beta_{1}=0.9,\,x_{0}^{1}=3,\,x_{0}^{2}=2$. The weighted terminal wealth $\beta_{1}X_{1}^{*}$ of agent 1 is marked by a square in the respective color.
5.2. Two agent case with power utility
@@
where $\Phi$ denotes the cumulative distribution function of the standard normal distribution. Figures 8, 9(a) and 9(b) illustrate the Nash equilibrium in terms of $Z_{T}$ for the parameter choice $x_{0}^{1}=3$, $x_{0}^{2}=2$, $\gamma=0.7$ and $\alpha_{i},\beta_{i}$, $i=1,2,$ so that (3.20) holds.
Figure 8 shows the Nash equilibrium $(X_{1}^{*},X_{2}^{*})$ as a function of $Z_{T}$ for $\alpha_{2}=0.5$ and $\beta_{1}=0.9$. The purple and black solid line show the terminal wealth of agents 1 and 2 in the Nash equilibrium. For comparison, we also illustrated the unconstrained optimal terminal wealth of agent 2 as the green dashed line. Similar to the logarithmic case, the terminal wealth of agent 1 is again the solution to the standard problem without the VaR-based constraint and is thus continuous. The terminal wealth $X_{2}^{*}$ of agent 2 is smaller than the unconstrained terminal wealth for $Z_{T}<z_{1-\alpha_{2}}$ and larger for $Z_{T}\geq z_{1-\alpha_{2}}$ due to the structure of $X_{2}^{*}$ shown in Theorem 3.6. Note that due to $\gamma=0.7$ we are here in the setting where insured states are those with poor market performance. Figure 9(a) illustrates the influence of the parameter $\alpha_{2}$ on $X_{2}^{*}$. As expected, the location of the discontinuity of $X_{2}^{*}$ is decreasing in terms of $\alpha_{2}$ as it is simply given as the $(1-\alpha_{2})$-quantile of $Z_{T}$. Moreover, we notice that for $Z_{T}<z_{\alpha_{2}}$ the value of $X_{2}^{*}$ is decreasing in $\alpha_{2}$. This results from the budget constraint $\mathbb{E}[Z_{T}X_{2}^{*}]=x_{0}^{2}$ since for larger $\alpha_{2}$, the value of $X_{2}^{*}$ is kept on the larger value $\beta_{1}X_{1}^{*}$ on a larger interval. This leads to the effect that a high probability for the constraint in return implies a higher underperformance on uninsured sates.
Finally, Figure 9(b) shows the terminal wealth $X_{2}^{*}$ of agent 2 in the Nash equilibrium for different choices of the parameter $\beta_{1}$. We notice a change of order of the value of $X_{2}^{*}$ for the different choices of $\beta_{1}$ located at the discontinuity $z_{1-\alpha_{2}}.$ For $Z_{T}\geq z_{1-\alpha_{2}}$, the value of $X_{2}^{*}$ is largest for the largest choice of $\beta_{1}$ as the value of $X_{2}^{*}$ is kept at $\beta_{1}X_{1}^{*}$ in this case. For $Z_{T}<z_{1-\alpha_{2}}$, the order of the values is opposite, i.e. the largest choice of $\beta_{1}$ yields the smallest value of $X_{2}^{*}$. This is again due to the budget constraint $\mathbb{E}[Z_{T}X_{2}^{*}]=x_{0}^{2}$. As in the logarithmic case the parameter $\beta_{1}$ has no influence on the set of insured states.
@@
5.3. More than two agents with logarithmic utility
It remains to discuss Nash equilibria for more than two agents using logarithmic utilities. Here, we restrict to the case that the sum $\alpha_{1}+\ldots+\alpha_{n}$ is less or equal to $1$. Thus, we can use Theorem 4.2 to compute the terminal wealth of the agents in the Nash equilibrium. Figure 10 displays the terminal wealths of $n=4$ agents in the Nash equilibrium for the parameter choice $\alpha_{i}=0.2$, $\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}=0.3$ for all $i,j\in\{1,\ldots,4\},i\neq j$ and $x_{0}^{1}=5$, $x_{0}^{2}=4$, $x_{0}^{3}=3$, $x_{0}^{4}=2$. The sets $A_{i}$, $i=1,\ldots,4,$ are chosen as $A_{i}=(z_{\alpha_{1}+\ldots+\alpha_{i-1}},z_{\alpha_{1}+\ldots+\alpha_{i}}]$, $i=1,\ldots,4$, where $z_{\alpha_{1}+\ldots+\alpha_{i}}:=0$ for $i=0$. We notice that the terminal wealth of agent 4 is larger than the terminal wealth of agent 3 on the set $A_{4}$ although agent 4 starts with a smaller initial capital. However, this results in a terminal wealth on $A_{4}^{c}$ that is significantly smaller than the optimal terminal wealth in the respective unconstrained problems. A comparison of $X_{3}^{*}$ and $X_{4}^{*}$ to the solutions of the respective unconstrained problems can be found in Figures 11(a) and 11(b). To sum up, we obviously see that wealthy investors are not or only marginally affected by the competition, whereas poorer agents deviate dramatically from their benchmark optimal strategy, leading to very volatile and risky investments at least in some scenarios. Moreover, in order to obtain a Nash equilibrium, the agents have to communicate and identify their sets where they want to ensure the target.
Figure 10. Nash equilibrium $(X_{1}^{*},X_{2}^{*},X_{3}^{*},X_{4}^{*})$ from Theorem 4.2 in terms of $Z_{T}$ for $\alpha_{i}=0.2$, $\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}=0.3$ for $i\in\{1,\ldots,4\},i\neq j$, $x_{0}^{1}=5$, $x_{0}^{2}=4$, $x_{0}^{3}=3$, $x_{0}^{4}=2$ and $A_{i}=(z_{\alpha_{1}+\ldots+\alpha_{i-1}},z_{\alpha_{1}+\ldots+\alpha_{i}}]$, $i=1,\ldots,4$, where $z_{\alpha_{1}+\ldots+\alpha_{i}}:=0$ for $i=0$.
@@ 6. Conclusion
It is meaningful to consider competing agents by introducing probability constraints instead of relative wealth targets, because this is in line with observations from psychologists and resembles incentives for fund managers. However, it turns out that finding Nash equilibria is in general more challenging. An interesting case which is analytically solvable is the case that all investors have a logarithmic utility function. In this case, Nash equilibria (if they exist) can be found for an arbitrary number of agents. In order to succeed, a number of parameter cases have to be distinguished leading either to no Nash equilibrium, a unique Nash equilibrium or an infinite number of Nash equilibria. We illustrate our findings by numerical examples. It turns out that wealthy investors are not or only marginally affected by the competition, whereas poorer agents deviate dramatically from their benchmark optimal strategy, leading to very volatile and risky investments at least in some scenarios. For other utility functions, and in particular, if the agents have different utility functions or also different information, the problem is much more demanding and we leave it for future research.
@@ 7.3. Maximizing L
$L(X,\lambda_{2},\eta_{2}):=U(X)-\lambda_{2}Z_{T}X+\eta_{2}\mathbbm{1}_{[X\geq\beta_{1}I(\kappa Z_{T})]}$
with $\lambda_{2},\eta_{2}\geq 0$ fixed. We show that $L$ is maximized by $X_{2}^{*}$ from (3.22). Obviously the maximum points can either be $I(\lambda_{2}Z_{T})$ or $\beta_{1}I(\kappa Z_{T})$ where $I(\lambda_{2}Z_{T})<\beta_{1}I(\kappa Z_{T}).$ We can compare the two possible values of the function $L$:
$F(Z_{T}):=U(I(\lambda_{2}Z_{T}))-\lambda_{2}Z_{T}I(\lambda_{2}Z_{T})-\Big{\{}U(\beta_{1}I(\kappa Z_{T}))-\lambda_{2}Z_{T}\beta_{1}I(\kappa Z_{T})+\eta_{2}\Big{\}}.$
@@
$F^{\prime}(Z_{T})$ $=\lambda_{2}\big{(}\beta_{1}I(\kappa Z_{T})-I(\lambda_{2}Z_{T})\big{)}-\beta_{1}\kappa Z_{T}I^{\prime}(\kappa Z_{T})(\kappa\beta_{1}^{-\gamma}-\lambda_{2})$
$=Z_{T}^{-1/\gamma}\bigg{(}\lambda_{2}\Big{(}\beta_{1}\kappa^{-1/\gamma}-\lambda_{2}^{-1/\gamma}\Big{)}+\frac{\beta_{1}\kappa^{-1/\gamma}}{\gamma}\Big{(}\kappa\beta_{1}^{-\gamma}-\lambda_{2}\Big{)}\bigg{)}=:Z_{T}^{-1/\gamma}f(\lambda_{2}).$
@@
$f\big{(}\kappa\beta_{1}^{-\gamma}\big{)}=0.$
@@
$f^{\prime}(\lambda_{2})$ $=\Big{(}1-1/\gamma\Big{)}\Big{(}\beta_{1}\kappa^{-1/\gamma}-\lambda_{2}^{-1/\gamma}\Big{)},$
so that $f^{\prime}(\lambda_{2})<0$ for all $\lambda_{2}>\kappa\beta_{1}^{-\gamma}$ (in case $10<\gamma<$ 1. The inequality reverses if $\gamma>1$). Thus, since we already saw in the proof of Theorem 3.6 that $\lambda_{2}>\kappa\beta_{1}^{-\gamma}$, we deduce that $F^{\prime}(Z_{T})<0$ for all values of $Z_{T}$. Moreover, by definition of $\eta_{2}$, we have $F(z_{1-\alpha_{2}})=0$ which implies that $X_{2}^{*}$ maximizes the function $L$.
7.4. Replicating the terminal wealth in the Nash equilibrium
In the following, we provide an auxiliary statement used to find the replicating strategies for the Nash equilibria found in Theorem 3.3 in the special case of a Black-Scholes market. Thus, we use the financial market explained at the beginning of Section 2. Additionally, we introduce the state price density process given by
$Z_{t}=\exp\Big{(}-\theta^{\top}W_{t}-\frac{1}{2}\|\theta\|^{2}t\Big{)},\quad 0\leq t\leq T.$
Then a straightforward application of Theorem E.1 in Jin and Yu Zhou, (2008) yields the following result, where $\varphi$ and $\Phi$ denote the density and cumulative distribution function of the standard normal distribution.
Lemma 7.3.
Let $0\leq c_{1}<c_{2}\leq\infty$ and $X=Z_{T}^{-1}\mathbbm{1}\{c_{1}<Z_{T}<c_{2}\}$. If $0<c_{1}<c_{2}<\infty$, the wealth-portfolio pair replicating $X$ is given by
$X(t)$ $=\frac{1}{Z_{t}}\Bigg{(}\Phi\Bigg{(}\frac{\log(c_{2})-\log(Z_{t})+\frac{1}{2}\|\theta\|^{2}(T-t)}{\|\theta\|\sqrt{T-t}}\Bigg{)}-\Phi\Bigg{(}\frac{\log(c_{1})-\log(Z_{t})+\frac{1}{2}\|\theta\|^{2}(T-t)}{\|\theta\|\sqrt{T-t}}\Bigg{)}\Bigg{)},$
$\pi(t)$ $=\Bigg{[}X(t)+\frac{1}{Z_{t}\|\theta\|\sqrt{T-t}}\Bigg{(}\varphi\Bigg{(}\frac{\log(c_{2})-\log(Z_{t})+\frac{1}{2}\|\theta\|^{2}(T-t)}{\|\theta\|\sqrt{T-t}}\Bigg{)}$
$\hskip 128.0374pt-\varphi\Bigg{(}\frac{\log(c_{1})-\log(Z_{t})+\frac{1}{2}\|\theta\|^{2}(T-t)}{\|\theta\|\sqrt{T-t}}\Bigg{)}\Bigg{)}\Bigg{]}\big{(}\sigma\sigma^{\top}\big{)}^{-1}\mu.$
The portfolio strategy $\pi$ describes the amount of money invested into the $d$ stocks at each time $t\in[0,T)$. Results for the cases $c_{1}=0$ and $c_{2}=\infty$ can be obtained by taking the respective limits.
Proof.
The result is a straightforward application of Theorem E.1 in Jin and Yu Zhou, (2008) using the following auxiliary calculation:
$\int_{a}^{b}\frac{1}{y}\varphi\Big{(}\frac{\log(y)-\nu}{\tau}\Big{)}\mathrm{d}y$ $=\tau\int_{a}^{b}\underbrace{\frac{1}{\tau y}\varphi\Big{(}\frac{\log(y)-\nu}{\tau}\Big{)}}_{(*)}\mathrm{d}y$
$=\tau\left(\Phi\Big{(}\frac{\log(b)-\nu}{\tau}\Big{)}-\Phi\Big{(}\frac{\log(a)-\nu}{\tau}\Big{)}\right),$
where we used that ()(*)( ∗ ) is the density function of a lognormal distribution with parameters $\nu$ and $\tau^{2}$ and that the associated cumulative distribution function is given by $\Phi((\log(\cdot)-\nu)/\tau)$. ∎